Mike is 6 years younger than Cassie. Twelve years ago, CCassie was twice as old as Mike was. How old is each now?
Also, Shalane is 14 and her grandmother is 54. How many years ago was her grandmother 6 times as old as Shalane?
If you could just help a bit, that would be great. Thankyou

Answer by stanbon(75887)

Mike is 6 years younger than Cassie.
Twelve years ago, Cassie was twice as old as Mike was. How old is each now?
----
m = c-6
(c-12) = 2(m-12)
---
Substitute for "m" and solve for "c":
c-12 = 2(c-6-12)
c-12 = 2(c-18)
c-12 = 2c-36
c = 24 (Cassie's age now)
m = c-6 = 18 (Mikes age now)
=================================

Also, Shalane is 14 and her grandmother is 54. How many years ago was her grandmother 6 times as old as Shalane?
Shalane = 14
Grandma = 54
----
Equation:
(54-x) = 6(14-x)
54-x = 84 - 6x
5x = 30
x = 6 (# of years ago)
==========================
Cheers,
Stan H.
=============